OpenStudy (anonymous):

Using the limit comparison test determine if the following converges or diverges

OpenStudy (anonymous):

\[\sum_{\infty}^{n=1}\tan \frac{ 1 }{ n }\]

OpenStudy (zarkon):

compute \(\displaystyle\Large\lim_{n\to\infty}\frac{\tan(\frac{1}{n})}{\frac{1}{n}}\)